# SlipStream: Adapting Pipelines for Distributed Training of Large DNNs Amid Failures

### v1

@@ Abstract.
Training large Deep Neural Network (DNN) models requires thousands of GPUs for days or weeks at a time. At these scales, failures are frequent and can have a big impact on training throughput. Restoring performance using spare GPU servers becomes increasingly expensive as models grow.
SlipStream is a system for efficient DNN training in the presence of failures, without using spare servers. It exploits the functional redundancy inherent in distributed training systems – servers hold the same model parameters across data-parallel groups – as well as the bubbles in the pipeline schedule within each data-parallel group. SlipStream dynamically re-routes the work of a failed server to its data-parallel peers, ensuring continuous training despite multiple failures. However, re-routing work leads to imbalances across pipeline stages that degrades training throughput. SlipStream introduces two optimizations that allow re-routed work to execute within bubbles of the original pipeline schedule. First, it decouples the backward pass computation into two phases. Second, it staggers the execution of the optimizer step across pipeline stages. Combined, these optimizations enable schedules that minimize or even eliminate training throughput degradation during failures. We describe a prototype for SlipStream and show that it achieves high training throughput under multiple failures, outperforming recent proposals for fault-tolerant training such as Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 ×, respectively.
copyright: none
@@ 1. Introduction
Two recent projects have exploited pipeline parallelism instead. Bamboo (Thorpe et al., 2023) relies on redundant computation and executes each pipeline stage on two nodes even in the fault-free case. When a node fails, the alternative nodes for its stages will do the work. Bamboo’s drawback is the drop in fault-free training throughput due to the redundant computations and added memory pressure on each GPU node. Oobleck (Jang et al., 2023) avoids overheads in fault-free execution by pre-computing a number of templates for pipeline parallelism that use a different number of nodes. As nodes fail, Oobleck replaces the original pipeline template with one that uses fewer nodes. The performance challenge for Oobleck is balancing work across heterogeneous pipelines as the pipelines with fewer nodes can become stragglers. Both Bamboo and Oobleck require significant re-configuration as nodes fail or rejoin, which can be a challenge as failure rates increase.
@@
In this paper, we introduce SlipStream, a novel scheme for resilient training with hybrid-parallel systems. SlipStream enables efficient execution in the presence of failures, without the need for spares and without any impact on model accuracy. Similar to Bamboo and Oobleck, we exploit pipeline parallelism. But unlike these previous proposals that handle failures within a single pipeline, SlipStream utilizes the inherent functional redundancy across pipelines in hybrid-parallel training systems. SlipStream re-routes the micro-batches of failed nodes to peer nodes that process the same pipeline stage for the model in the other data-parallel groups. Peer nodes store the same model parameters and can handle the extra work without the need for parameter re-shuffling.
@@
If done naively, SlipStream’s adaptive pipelining increases training time and memory usage due to the additional work for the peers of failed nodes. SlipStream uses a series of optimizations that exploit unique characteristics of hybrid-parallel training to eliminate all or most of the inefficiencies. SlipStream uses the bubbles in the pipeline schedule of peers to execute the additional work at low or no overhead. These bubbles typically appear during the start-up and cool-down phase of the pipeline schedule, but not during the steady-state when most additional work must occur. SlipStream overcomes this challenge by decoupling back propagation for micro-batches into two distinct gradient computations, one relative to the input and one relative to the parameters. The two steps are independently scheduled in order to better leverage bubbles in the the cool-down portion of the pipeline schedule. The selective application of decoupled back propagation exploits the imbalance in memory usage across stages in hybrid-parallel training while avoiding an increase in peak memory pressure. Finally, SlipStream staggers the execution of the optimizer tasks across pipeline stages in order to leverage bubbles from the start-up phase of the pipeline schedule for the next training iteration.
@@
We implemented SlipStream on top of the DeepSpeed training framework (Rasley et al., 2020). The key SlipStream element is the Planner that uses dynamic programming (DP) and mixed integer linear programming (MILP) to find the best pipeline schedule in the presence of multiple failures. The Planner runs offline and calculates efficient schedules for up to a certain number of failures. These schedules are employed as needed during training. We evaluated SlipStream using DNNs with billions of parameters like GPT-3 (Brown et al., 2020), using both real-world experiments and simulations of large training systems. We show that SlipStream supports high throughput training even at high failure counts, e.g. $\geq$10of the overall system. SlipStream outperforms Oobleck by 1.46×1.46$\times$1.46 × in training scenarios with realistic GPU failures. SlipStream’s advantage stems from its efficient scheduling of re-routed work and its ability to avoid major re-shuffling of model parameters upon failures. SlipStream outperforms Bamboo by 1.64×1.64$\times$1.64 ×. It is also able to efficiently train much larger models than Bamboo, which requires large memory overheads for its fault tolerance mechanism.
@@ 2.1. Distributed DNN Training
Large-scale training systems balance these tradeoffs by utilizing all three forms of parallelism as shown in Figure 1 (Narayanan et al., 2021; Li and Hoefler, 2021; Liu et al., 2023; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022). This is known as hybrid-parallelism. In the most common case, systems such as DeepSpeed (Rasley et al., 2020) or Alpa (Zheng et al., 2022) use a combination of tensor parallelism within a multi-GPU server, pipeline parallelism across multi-GPU servers, and data parallelism across pipelines. Hybrid parallelism enables large model training with graceful scaling and reasonable training times (weeks to few months) using optimized clusters with 1000s of densely-connected GPUs.
@@ 2.2. Fault Tolerance in Distributed DNN Training
Enhancing fault tolerance in distributed training involves three key issues: fast error detection, efficient checkpointing, and robust execution under faults. Fast error detection is crucial for minimizing downtime and preventing the spread of errors. Some errors can be quickly detected through encoding and redundancy mechanisms in the hardware (e.g., error correction codes) or software timeouts. Other errors are detected if they eventually lead to irregular spikes in the loss function optimized during training (Beaumont, 2022). The latter category includes silent data corruptions (Bacon, 2022; Dattatraya Dixit et al., 2022). Checkpointing, which saves the state of the model at regular intervals, allows for recovery from the recent stable state without restarting the entire training from scratch. Since naive checkpointing to remote storage leads to long pauses in the training process, there has been considerable research in improving its efficiency (Mohan et al., 2021; Eisenman et al., 2022; Wang et al., 2023; Jiang et al., 2024) as discussed in Section 6.
2.3. Efficient DNN Training in Presence of Failures
@@
Our work focuses on robust training in the presence of faults. A common approach is to have a reserve of spares ready to replace servers with failed GPU servers. To reduce recovery time, spare servers can also be hot standbys with their memory loaded with model parameters at the end of each iteration. The key advantage of spare servers is that, upon recovery, training efficiency returns to its pre-failure state. However, this approach becomes increasingly costly as systems scale up, fault frequency increases, and more spares are needed. Spare servers must connect to the network of the training cluster with a similar bandwidth profile as the servers they will be called to replace, which incurs additional networking costs and typically requires that spares are allocated in coarse-grained groups (Jouppi et al., 2023). Hot standby servers should be provisioned according to the full capacity of a data parallel group, determined by multiplying pipeline parallelism ($PP$) by tensor parallelism ($TP$). For instance, training a 530B GPT model requires 280 spare GPUs on standby, which results in an 11% increase in training costs (Narayanan et al., 2021). Spare servers can potentially run small training jobs while on (non hot) standby to amortize cost. However, the job eviction latency needed to free these resources upon a failure can cause long stalls for large training jobs occupying thousands of GPUs.
The alternative approach is to continue training with the largest possible subset of functional resources. Ideally, when $x\%$ of GPU servers fail, training can continue at no less than $(100-x)\%$ of the fault-free throughput. The simplest method, sometimes referred to as elastic batching, is to drop a whole data parallel (DP) group when a node fails (Wang et al., 2018; Lin et al., 2019). The remaining data parallel groups can proceed with the same parallelization scheme. While straightforward, this method has substantial drawbacks. A single node failure results in $PP\times TP$ nodes going offline, decreasing the total throughput by $1/DP$ (e.g., 17% when training a 1000B parameter GPT model (Narayanan et al., 2021)). Moreover, this requires reducing the training batch size or recompiling the program to optimize each data parallel group for a larger portion of the batch. This can disrupt the balance and efficiency of the entire training job.
Figure 2. Adaptive pipelining when $W_{1\textunderscore 2}$ fails. The output from worker $W_{1\textunderscore 1}$ is redirected to workers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$.
@@
Figure 3. Hybrid-parallel training with 3 data-parallel pipelines, 4 pipeline stages, and 6 micro-batches per iteration.  3(a) shows a 1F1B fault-free schedule.  3(b) shows how Adaptive Pipelining re-routes micro-batches failed workers to functional peers.
@@
Two recent projects have instead utilized pipeline parallelism for efficient training in the presence of failures without spares. Bamboo (Thorpe et al., 2023) draws inspiration from RAID (Patterson et al., 1988) and designs a pipeline for redundant computation (RC). Each pipeline stage is redundantly computed on two nodes, even in the fault-free case. When a node fails, its backup node steps up and computes the forward and backward passes of the failed node. Bamboo uses RC even in the absence of faults. While RC tries to hide some overhead in pipeline bubbles, its extra computational overhead still significantly drops training throughput (see Section 5). RC also replicates the state of each pipeline stage in two nodes, increasing GPU memory pressure and limiting its scalability. Finally, Bamboo needs to restart with a full reconfiguration from a checkpoint for as few as two failures, e.g., when two adjacent nodes fail.
Oobleck (Jang et al., 2023) adapts pipeline parallelism for resilient execution but has no overhead in the fault-free case. It relies on pre-computed templates for pipeline parallelism. All templates are logically equivalent yet physically heterogeneous. Each template has a different number of stages, a different micro-batch configuration, and occupies a different number of nodes. A pipeline with failed node(s) is replaced with a new pipeline from a template that requires fewer nodes. However, templates with fewer nodes have lower training throughput. To avoid having the slowest pipeline slowdown the overall synchronous training job, Oobleck attempts to distribute the global mini-batch proportionally to compute power. This is not always possible and the straggling pipelines will inflate training time. Moreover, each pipeline re-configuration is in the critical path of training and can become an overhead if nodes fail frequently. Re-configuration is also needed when nodes re-join. Finally, Oobleck treats each heterogeneous pipeline as a black-box. It fails to observe that there are bubbles across all pipelines that can help hide some or all of the overheads due to node failures.
3. SlipStream Techniques
@@
SlipStream aims to support efficient distributed training in the presence of faults. It requires no spare servers and has no impact on model accuracy compared to fault-free training. SlipStream can tolerate multiple hardware failures and maintains training throughput proportional to the number of functional servers available. SlipStream is optimized for fast recovery as failures do not require significant re-shuffling of model parameters between functional nodes.
@@
This section reviews the key SlipStream techniques: Adaptive Pipelining, Decoupled BackProp, and a Staggered Optimizer. Section 4 presents the SlipStream system design.
@@ 3.1. Adaptive Pipelining: Working Around Failures
SlipStream exploits two key properties in hybrid-parallel training systems. First, there is functional redundancy across data-parallel pipelines. The GPUs that process the same pipeline stage hold identical parameters and only differ in the micro-batches they process. In Figure 2, for example, workers $W_{0\_2}$, $W_{1\_2}$, and $W_{2\_2}$ are peers that hold identical parameters for stage 2 of the 4-stage pipeline. Second, there are bubbles (idle slots) in the pipeline schedule for each worker. The 1F1B pipeline schedule (Narayanan et al., 2019) in Figure 3(a) has 9 bubbles in the repeating 27-slot schedule for worker $W_{0\_2}$.
Adaptive Pipelining exploits functional redundancy and bubbles by dynamically re-routing micro-batches from a failed worker to its functioning peers in other data-parallel pipelines. We aim to use the bubbles in peers’ schedules to process the micro-batches for the failed worker at low performance penalty. We evenly distribute micro-batches across all functional peers. Since all pipelines perform a synchronized all-reduce at the end of each iteration, load balancing ensures that no single worker (and thus pipeline) is overloaded, delaying the progress of the entire training iteration.

### v2

# ReCycle: Resilient Training of Large DNNs using Pipeline Adaptation

@@ Abstract.
Training large Deep Neural Network (DNN) models requires thousands of GPUs over the course of several days or weeks. At this scale, failures are frequent and can have a big impact on training throughput. Utilizing spare GPU servers to mitigate performance loss becomes increasingly costly as model sizes grow. ReCycle is a system designed for efficient DNN training in the presence of failures, without relying on spare servers. It exploits the inherent functional redundancy in distributed training systems – where servers across data-parallel groups store the same model parameters – and pipeline schedule bubbles within each data-parallel group. When servers fails, ReCycle dynamically re-routes micro-batches to data-parallel peers, allowing for uninterrupted training despite multiple failures. However, this re-routing can create imbalances across pipeline stages, leading to reduced training throughput. To address this, ReCycle introduces two key optimizations that ensure re-routed micro-batches are processed within the original pipeline schedule’s bubbles. First, it decouples the backward pass into two phases: one for computing gradients for the input and another for calculating gradients for the parameters. Second, it avoids synchronization across pipeline stages by staggering the optimizer step. Together, these optimizations enable adaptive pipeline schedules that minimize or even eliminate training throughput degradation during failures. We describe a prototype for ReCycle and show that it achieves high training throughput under multiple failures, outperforming recent proposals for fault-tolerant training such as Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 ×, respectively.
Fault-tolerant Training, Distributed Training, Hybrid Parallelism, Pipeline Adaptation
ccs: Computer systems organization Dependable and fault-tolerant systems and networksccs: Computer systems organization Distributed architecturesccs: Computer systems organization Neural networkscopyright: acmlicensedjournalyear: 2024conference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695960isbn: 979-8-4007-1251-7/24/11
@@ 1. Introduction
Two recent projects have exploited pipeline parallelism instead. Bamboo (Thorpe et al., 2023) relies on redundant computation and executes each pipeline stage on two nodes even in the fault-free case. When a node fails, the alternative nodes for its stages will do the work. Bamboo’s drawback is the drop in fault-free training throughput due to the redundant computations and added memory pressure on each GPU node. Oobleck (Jang et al., 2023) avoids overheads in fault-free execution by precomputing a number of templates for pipeline parallelism that use a different number of nodes. As nodes fail, Oobleck replaces the original pipeline template with one that uses fewer nodes. The performance challenge for Oobleck is balancing work across heterogeneous pipelines as the pipelines with fewer nodes can become stragglers. Both Bamboo and Oobleck require significant re-configuration as nodes fail or rejoin, which can be a challenge as failure rates increase.
@@
In this paper, we introduce ReCycle, a novel scheme for resilient training with hybrid-parallel systems. ReCycle enables efficient execution in the presence of failures, without the need for spares and without any impact on model accuracy. Similar to Bamboo and Oobleck, we exploit pipeline parallelism. But unlike these previous proposals that handle failures within a single pipeline, ReCycle utilizes the inherent functional redundancy across pipelines in hybrid-parallel training systems. ReCycle re-routes the micro-batches of failed nodes to peer nodes that process the same pipeline stage for the model in the other data-parallel groups. Peer nodes store the same model parameters and can handle the extra work without the need for parameter re-shuffling.
@@
If done naively, ReCycle’s adaptive pipelining increases training time and memory usage due to the additional work for the peers of failed nodes. ReCycle uses a series of optimizations that exploit unique characteristics of hybrid-parallel training to eliminate all or most of the inefficiencies. ReCycle uses the bubbles in the pipeline schedule of peers to execute the additional work at low or no overhead. These bubbles typically appear during the start-up and cool-down phase of the pipeline schedule, but not during the steady-state when most additional work must occur. ReCycle overcomes this challenge by decoupling back propagation for micro-batches into two distinct gradient computations, one relative to the input and one relative to the parameters. The two steps are independently scheduled in order to better leverage bubbles in the the cool-down portion of the pipeline schedule. The selective application of decoupled back propagation exploits the imbalance in memory usage across stages in hybrid-parallel training while avoiding an increase in peak memory pressure. Finally, ReCycle staggers the execution of the optimizer tasks across pipeline stages in order to leverage bubbles from the start-up phase of the pipeline schedule for the next training iteration.
@@
We implemented ReCycle on top of the DeepSpeed training framework (Rasley et al., 2020). The key component of ReCycle is the Planner, which utilizes dynamic programming and mixed-integer linear programming (MILP) to determine adaptive pipeline schedule in presence of multiple failures. Operating offline, the Planner precomputes efficient schedules for a predefined number of failures, which are then applied as necessary during training. We evaluated ReCycle using DNNs with billions of parameters like GPT-3 (Brown et al., 2020), using both real-world experiments and simulations of large training systems. We show that ReCycle supports high throughput training even at high failure counts, e.g. $\geq$10of the overall system. ReCycle outperforms Oobleck by 1.46×1.46$\times$1.46 × in training scenarios with realistic GPU failures. ReCycle’s advantage stems from its efficient scheduling of re-routed work and its ability to avoid major re-shuffling of model parameters upon failures. ReCycle outperforms Bamboo by 1.64×1.64$\times$1.64 ×. It is also able to efficiently train much larger models than Bamboo, which requires large memory overheads for its fault tolerance mechanism.
@@ 2.1. Distributed DNN Training
Large-scale training systems balance these trade-offs by utilizing all three forms of parallelism as shown in Figure 1 (Narayanan et al., 2021; Li and Hoefler, 2021; Liu et al., 2023; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022). This is known as hybrid-parallelism. In the most common case, systems such as DeepSpeed (Rasley et al., 2020) or Alpa (Zheng et al., 2022) use a combination of tensor parallelism within a multi-GPU server, pipeline parallelism across multi-GPU servers, and data parallelism across pipelines. Hybrid parallelism enables large model training with graceful scaling and reasonable training times (weeks to few months) using optimized clusters with 1000s of densely-connected GPUs.
@@ 2.2. Fault Tolerance in Distributed DNN Training
Enhancing fault tolerance in distributed training requires addressing three critical areas: fast error detection, checkpointing, and efficient execution in presence of faults.
2.2.1. Error Detection
Fast Error Detection is vital for minimizing downtime and preventing the propagation of errors. Effective error detection mechanisms include both hardware and software solutions. Hardware mechanisms, such as error correction codes (ECC), provide immediate detection and correction of data corruption. Software solutions can involve timeouts and heartbeat signals to identify system failures quickly. Additionally, some errors, particularly those that are silent or not immediately apparent, are detected through irregularities in the loss function. These anomalies, often indicative of silent data corruptions, can significantly impact training outcomes if not addressed promptly (Beaumont, 2022; Bacon, 2022; Dattatraya Dixit et al., 2022).
2.2.2. Checkpointing
Checkpointing plays a crucial role in fault tolerance by allowing the system to recover from a recent stable state rather than restarting the entire training process. It involves saving the model’s state at regular intervals. However, naive checkpointing methods that write data to remote storage can introduce significant pauses in the training process due to latency and bandwidth constraints. To address this, extensive research has focused on optimizing checkpointing techniques to reduce overhead and improve efficiency (Mohan et al., 2021; Eisenman et al., 2022; Wang et al., 2023; Jiang et al., 2024), as detailed in Section 7.
2.2.3. Efficient DNN Training in the Presence of Failures
@@
Our work focuses on ensuring that the system remains operational even when parts of it fail. A common method is to use warm spares, preloaded with model parameters to reduce recovery time. While this ensures that training efficiency returns to pre-failure levels, it becomes costly as systems scale and fault frequency rises. Spare servers need similar network bandwidth as the ones they replace, adding networking costs and requiring coarse-grained allocation (Jouppi et al., 2023). For example, training a 530B GPT model requires 280 spare GPUs, provisioned according to the full capacity of a data parallel group, increasing costs by 11% (Narayanan et al., 2021). Use of cold spares to run small jobs can offset some of these costs, however eviction delays upon failure can cause significant stalls for large jobs.
Figure 2. Adaptive pipelining when $W_{1\textunderscore 2}$ fails. The micro-batches from worker $W_{1\textunderscore 1}$, originally intended for $W_{1\textunderscore 2}$, are dynamically re-routed to workers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$, ensuring that the training process continues without interruption.
@@
Figure 3. Hybrid-parallel training across 12 workers with 3 data-parallel pipelines, each with 4 pipeline stages, and 6 micro batches.  3(a) shows a 1F1B fault-free schedule.  3(b) shows how Adaptive Pipelining re-routes micro-batches from failed worker $W_{1\textunderscore 2}$ to its functional peers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$.
An alternative approach is to continue training with the largest subset of available resources. Ideally, if $x\%$ of GPU servers fail, training proceeds at no less than $(100-x)\%$ of fault-free throughput. A simple method, known as elastic batching, drops an entire data parallel group when a node fails, allowing the remaining groups to continue with the same parallelization (Wang et al., 2018; Lin et al., 2019). However, this approach has significant drawbacks, as a single node failure can take $PP\times TP$ nodes offline, reducing throughput by $1/DP$ (e.g., 17% for a 1T parameter GPT model (Narayanan et al., 2021)). Moreover, this approach requires reducing the batch size or recompiling the program to optimize each data parallel group for a larger batch portion, disrupting the balance and efficiency of the training job.
@@
Two recent projects use pipeline parallelism for fault-tolerant training without spares. Bamboo (Thorpe et al., 2023), drawing inspiration from RAID (Patterson et al., 1988), employs redundant computation (RC) where each pipeline stage is replicated on two nodes, even in fault-free cases. When a node fails, its backup handles its and forward and backward passes of failed node. Though RC hides some overhead in pipeline bubbles, it still significantly reduces throughput (see Section 6) and increases GPU memory pressure, limiting scalability. Moreover, Bamboo must restart with a full reconfiguration from a checkpoint for as few as two adjacent node failures.
Oobleck (Jang et al., 2023) uses pipeline parallelism to ensure resilient execution with no overhead when no faults occur. It utilizes precomputed templates for pipeline parallelism, each differing in stages, micro-batch configurations, and node counts. If a pipeline encounters a failed node, it switches to a template with fewer nodes, which may reduce training throughput. To prevent a slow pipeline from affecting the overall job, Oobleck distributes the global mini-batch based on compute power, leading to increased training time. Pipeline re-configuration when nodes fail or re-join can also become an overhead if failures are frequent. Additionally, Oobleck treats each pipeline as a black-box and does not leverage bubbles to mitigate overheads from node failures.
3. ReCycle Techniques
@@
ReCycle aims to support efficient distributed training in the presence of faults. It requires no spare servers and has no impact on model accuracy compared to fault-free training. ReCycle can tolerate multiple hardware failures and maintains training throughput proportional to the number of functional servers available. ReCycle is optimized for fast recovery as failures do not require significant re-shuffling of model parameters between functional nodes.
@@
This section reviews the key ReCycle techniques: Adaptive Pipelining, Decoupled BackProp, and a Staggered Optimizer. Section 4 presents the ReCycle system design.
@@ 3.1. Adaptive Pipelining: Working Around Failures
ReCycle exploits two key properties in hybrid-parallel training systems. First, there is functional redundancy across data-parallel pipelines. The GPUs that process the same pipeline stage hold identical parameters and only differ in the micro-batches they process. In Figure 2, for example, workers $W_{0\_2}$, $W_{1\_2}$, and $W_{2\_2}$ are peers that hold identical parameters for stage 2 of the 4-stage pipeline. Second, there are bubbles (idle slots) in the pipeline schedule for each worker. The 1F1B pipeline schedule (Narayanan et al., 2019) in Figure 3(a) has 9 bubbles in the repeating 27-slot schedule for worker $W_{0\_2}$.
Adaptive Pipelining exploits functional redundancy and bubbles by dynamically re-routing micro-batches from a failed worker to its functioning peers in other data-parallel pipelines. We aim to use the bubbles in peers’ schedules to process the micro-batches for the failed worker with a low performance penalty. We evenly distribute micro-batches across all functional peers. Since all pipelines perform a synchronized all-reduce at the end of each iteration, load balancing ensures that no single worker (and thus pipeline) is overloaded, delaying the progress of the entire training iteration.
@@
Are sufficient bubbles available? Adaptive Pipelining exploits the $3\times(PP-1)\times DP$ existing but previously unused bubbles in each pipeline to recover from failures. For example, the 405 billion parameter LLaMA-3 is trained with synchronous 1F1B over $8192$ GPUs using hybrid parallelism with tensor, pipeline, and data parallelism degree $TP=8$, $PP=16$, and $DP=64$, respectively (AI, 2024). ReCycle can leverage $(3\times(16-1)\times 64)=2880$ bubbles per iteration to accommodate $\frac{2880}{3}=960$ rerouted micro-batches per iteration. For a training job with global batch size of 2048, this is sufficient to handle 30 simultaneous failures.
The scheduling challenge: While sufficient bubbles exist, these bubbles are concentrated in the warm-up and cool-down phase of the 1F1B schedule. For example, there are 18 idle slots across data-parallel peers of failed node $W_{1\_2}$ in Figure 3(a) to accommodate its micro-batches. Adaptive Pipelining needs to reroute micro-batches to peer workers $W_{0\_2}$ and $W_{2\_2}$ mostly in the middle of their steady-state schedule, which is optimized to be bubble-free for efficiency (Narayanan et al., 2019). As shown in Figure 3(b), this has a significant performance impact. The additional micro-batches to peer workers and the dependencies between forward and backward pass computations lead to 9 additional time steps per iteration, for a total of 36. In other words, the failure of 8.3% of workers (1 out of 12) leads to a 33% slowdown in iteration time.
3.2. Decoupled BackProp: Filling Unused Bubbles